OpenStudy (anonymous):

PLEASE PLEASE HELP Use synthetic division to find the quotient and remainder when 6x^6-x^4+7x^2+6 is divided by x-2 Quotient:______ Remainder:_______

OpenStudy (zehanz):

Are you familiar with synthetic division?

OpenStudy (zehanz):

You would take the zero of x-2, that is 2, and all the coefficients of the polynomial, and order them like this: 2 | 6 0 -1 0 7 6 Remember, 6x^6-x^4+7x^2+6 is really \(6x^6+0x^5+-1x^4+0x^3+7x^2+6\).

OpenStudy (zehanz):

Now, the next step is, draw a line somewhat below the numbers, leaving some space between: 2 | 6 0 -1 0 7 6 | ----------------------- + Then just drop the first coefficient (6) below the line: 2 | 6 0 -1 0 7 6 | ----------------------- + 6

OpenStudy (zehanz):

Now, the actual work begins. Multiply 2 with the dropped 6 and put the result in the next column. Add: 2 | 6 0 -1 0 7 6 | 12 ----------------------- + 6 12

OpenStudy (zehanz):

Then, multiply 2 with the 12 below the line and put in the next column. Add: 2 | 6 0 -1 0 7 6 | 12 24 ----------------------- + 6 12 23 Keep doing this: 2 | 6 0 -1 0 7 6 | 12 24 46 92 198 ------------------------ + 6 12 23 46 99 204

OpenStudy (anonymous):

okay understand so far so do we keep using synthetic division?

OpenStudy (zehanz):

Awww, I see I still forgot one coefficient :( you have 0x, so we get: 2 | 6 0 -1 0 7 0 6 | 12 24 46 92 198 396 ----------------------------- + 6 12 23 46 99 198 402

OpenStudy (zehanz):

Now we are at the end. The first six numbers below the -------- line are the new coefficients of our 5th degree polynomial. The last (402) is the remainder. So: \[6x^6-x^4+7x^2+6=6x^5+12x^4+23x^3+46x^2+99x+198+\dfrac{402}{6x^6-x^4+7x^2+6}\]

OpenStudy (zehanz):

The end result is:\[6x^5+12x^4+23x^3+46x^2+99^x+198+\dfrac{402}{6x^6-x^4+7x^2+6}\]

OpenStudy (anonymous):

So the remainder is 402 and the quotient is 6x^5+12x^4+23x^3+46x^2+99x+98 ?? could it be more simplify

OpenStudy (zehanz):

Yes, the remainder and qouotient are as you said. It cannot be made simpler...

OpenStudy (anonymous):

thank you!!

OpenStudy (zehanz):

YW!